Evening shift 11/25/99

Richard Tesarek (tesarek@fnal.gov)
Fri, 26 Nov 1999 00:20:43 -0600

Shift Summary for evening shift, 25 Nov 99.

On Shift: RJT, SL

Hours of beam available at all intensities: 8.2
Hours of beam available at nominal intensity: 7.9
Hours of E799 data (special runs not included): 6.7
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: ktev2 crash (drt_fresh_start)
kfrends
pipeline reset
latch/DYC replacement

========================================================================
Runs ended this shift:
Run 15054 good data (ktev2 crash ended, no EOF on tapes)
Run 15055 good data, ktevana errors (latch problem)
Run 15056 debugging: data no good.

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:
Inherit relatively smooth running conditions from day shift. ktev2
crashes ending run 15054 and is rebooted. Attempt to save histograms
for this run, but l3#save hangs. Trigger 16 recovered toward end of
run 15054. drt_fresh_start after ktev2 reboot. Minor problems with
CsI (see below). Latch problems in trigger latch crate (crate 8).
Latch replaced and minor screwing around with cabling.

> ========================================================================
Description of Shift:



run spill time comment
------------------------------------------------------------------------
15054 16:15 Onsplit requests tape uph033 in ktev2/tps3d2.

107 16:23 Take advantage of empty spills to reseat cables
in trigger crate 11. Trigger 16 reappears this
spill.

16:55 Onsplit requests tape upt015 in ktev1/tps3d2.

17:21 ktev2 crashes. Attempt to stop run gracefully
fails. l3#save attempt to save histograms hangs.

15055 17:42 Start run

19:17 Onsplit requests tape upf048 in ktev4/tps3d2.

111-115 20:15 Val pipes up with "ped problem" and "laser
illumination". Discover 2 CsI channels (1941,1945)
drop from the laser illumination. Reset pipeline
and problem disappears by spill 115.

21:00 Onsplit requests tape upm036 in ktev1/tps3d2.

165- 21:45 ktevana errors, missing FFFF word from crate 8
(trigger latch crate). After numerous KFRENDs

stop run.

15056 21:51 Start run after replacing latch.
Trigger problems persist, replace DYC.

22:38 Discover problems with cabling after original
latch replacement. Stop run

15057 22:42 Start run. DC monitoring screen goes blank, FIXDISP.

========================================================================
Follow up Notes on Previous Shift

...none...
========================================================================
Follow up notes to next shift:

o Keep an eye out for problems with CsI channels listed below:
1096 - large gain drift (7%)
909 - wide stability rms (4.9%)
397 - wide pedestal rms (3.1%)

o drt_fresh_start at new run.
o HCC inefficient channels: 2168,2085,1533,1088,1089
o RC source timing for the 500mV threshold looks off for RC7-10
o RC7 500mV threshold may be wrong (more events than 200mV threshold).